OpenStudy (anonymous):

3xy^2+4y^2=4x^2-11x defines Y as a function of x.

OpenStudy (anonymous):

\[3xy^2+4y^2=4x^2-11x\]\[y^2(3x+4)=4x^2-11x\]\[y^2=\frac{4x^2-11x}{(3x+4)}\]\[y^2=\frac{x(4x-11)}{3x+4}\]\[y=\pm \sqrt{\frac{x(4x-11)}{3x+4}}\]